# Which symmetry group for elementary particles with an electric charge today and in the past?

### v1

@@ 8.3 Explicit form of the connection and the equation of motion
Our goal now is to determine explicitly the connection satisfying the hypotheses (H3) to (H5). As we know, the $\hat{\mathbb{G}}_{0}$-structure $\mathcal{F}_{\hat{\bm{G}_{0}}}$ is in general non integrable. Then, instead of the moving frame $(\hat{\bm{e}}^{\prime}_{1},\cdots,\hat{\bm{e}}^{\prime}_{5})$ of (H3), section of $\mathcal{F}_{\hat{\bm{G}_{0}}}$, we would like to work with a coordinate frame $(\bm{\partial}_{1},\cdots,\bm{\partial}_{4})$ associated to a coordinate system $(X^{1},\cdots,X^{4})$ of the space-time $\mathcal{U}$ and the corresponding fibered frame $(\hat{\bm{e}}_{1},\cdots,\hat{\bm{e}}_{5})$ adapted to the section $f$, such as $\hat{\bm{e}}_{\alpha}=(Tf)\bm{\partial}_{\alpha}$ and obtained by the change of basis
@@
As expected, according to the experience, the rest mass $m_{0}$ is an integral of the motion [^5] However, it is possible to extend the formalism to a particle with time-varying mass by introducing a thrust force as in de Saxcé and Vallée [2016]. Besides, (91) is reduced to
@@
represents a 1-covariant and 1-contravariant tensor $\bar{\bm{F}}$ of components $F^{i}_{j}$. The physical interpretation is that the vector $\bm{A}$ of which the adjoint $\bm{A}^{*}$ is represented by the row $A^{*}$ is the electromagnetic 4-potential, $\bm{F}$ (resp. $\bar{\bm{F}}$) is the 2-covariant (resp. 1 contravariant and 1-covariant) electromagnetic field 666with the conventions used in Souriau [1970, 1997b, §15] and
@@
Remark 3. it remains some freedom in the definition of the electromagnetic potential. For instance, we could have erased the factor $1/2$ before $A^{*}$ in the definition (79) but in this case we should have replaced $q$ by $-q/2$ everywhere. Both variants are conventional. We kept the one we considered the most convenient.
Remark 4. It is worth now to revisit the original Kaluza-Klein theory. In a $\hat{\bm{G}_{1}}$-basis $(\hat{\bm{e}}^{\prime}_{1},\cdots,\hat{\bm{e}}^{\prime}_{5})$, the metric is represented by
$\hat{G}^{\prime}=\left[\begin{array}[]{cc}G^{\prime}_{0}\hfill&0\hfill\\ 0\hfill&-1\hfill\\ \end{array}\right]$
As in general the Riemannian structures are non integrable, we consider a coordinate frame $(\bm{\partial}_{1},\cdots,\bm{\partial}_{5})$ associated to a coordinate system $(X^{1},\cdots,X^{5})$ of the 5 dimensional universe $\hat{\mathcal{U}}$ obtained by the change of basis
$\bm{\partial}_{\alpha}=\hat{P}^{\beta}_{\alpha}\hat{\bm{e}}^{\prime}_{\beta}$
with the transformation matrix (79). The metric is represented in the new frame by
$\hat{G}=\hat{P}^{T}\hat{G}^{\prime}\,\hat{P}=\left[{{\begin{array}[]{cc}G-4\,A^{*T}A^{*}\hfill&2\,A^{*T}\hfill\\ 2\,A^{*}\hfill&-1\hfill\\ \end{array}}}\right]$
where $G$ given by (80). Using classical tensor notation, it holds
$[\hat{G}_{\alpha\beta}]=\left[{{\begin{array}[]{cc}\hat{G}_{ij}=G_{ij}-4\,A_{i}A_{j}\hfill&\hat{G}_{j5}=2\,A_{j}\hfill\\ \hat{G}_{5i}=2\,A_{i}\hfill&\hat{G}_{55}=-1\hfill\\ \end{array}}}\right]$
If the expression stands out the standard presentation of the metric in Kaluza-Klein theory, it is matter of conventions. The factor $2$ before $A_{i}$ and $A_{j}$, and $-4$ before $A_{i}A_{j}$ result from the definition of $A^{*}$. We explained these things in the previous remark . The negative value of $\hat{G}_{55}$ is consistent with the choice of the signature (+)(+----)( + - - - - ). Restoring the positive value consists in adopting the signature (++++)(-++++)( - + + + + ).
Calculating the connection matrix, we have to take care that both $G$ and $A^{*}$ are fields defined on $\hat{\mathcal{U}}$, then functions of the five coordinates $\hat{X}^{1},\cdots,\hat{X}^{5}$, giving up Kaluza’s ”cylinder condition”.
Let $s\mapsto\hat{\bm{X}}(s)$ be the trajectory of a particle in the space-time, parameterized by the arc length $s$ (for the metric $\hat{\bm{G}}$) and
$\hat{\bm{U}}=\frac{d\hat{\bm{X}}}{ds}$
is the unit tangent vector such that $\hat{\bm{U}}^{*}\hat{\bm{U}}=1$. We claim that the motion of a particle is such that its linear 5-momentum is parallel-transported
^$\hat{\nabla}_{\hat{\bm{U}}}\hat{\bm{\Pi}}=\bm{0}$ (100)
The linear 5-momentum is represented in the $\hat{\bm{G}_{1}}$-basis by the 5-row
$\hat{\Pi}^{\prime}=[m_{0}\,U^{\prime*}P,q^{\prime}]$
with $q^{\prime}=m_{0}U^{\prime 5}$ and in the coordinate basis by
$\hat{\Pi}=[m_{0}\,U^{*}-2\,q\,A^{*},q]$
with $U^{*}=U^{\prime*}P$ and $q=m_{0}U^{5}$. Hence the trajectory of the particle is not a geodesic of the manifold.
@@ 8.4 Gauge transformation
The coordinate basis $(\bm{\partial}_{i})$ and $\hat{\bm{e}}_{5}$ being given, the adapted basis $(\hat{\bm{e}}_{\alpha})$ of $\mathcal{U}$ to $f$ is uniquely defined by the section $f$. We would like to know to what extend the equation of motion is independent on the choice of the section $f$. For this aim, we consider another section $f^{\prime\prime}$ and the corresponding adapted basis $(\hat{\bm{e}}^{\prime\prime}_{\alpha})$. As $(T\pi)\hat{\bm{e}}_{i}=(T\pi)\hat{\bm{e}}^{\prime\prime}_{i}=\bm{\partial}_{i}$, the vector $\hat{\bm{e}}^{\prime\prime}_{i}-\hat{\bm{e}}_{i}$ is tangent to the fiber of $\mathcal{U}$ over $\bm{X}$ then
@@ 9 Extended variational relativity
The goal of this section is to extend the general theorems of the relativity based on Einstein-Hilbert functional (Hilbert [1915]). We use Palatini formalism, considering the connection as an unknown independent field (Palatini [1919]).
@@ 9.1 Stationary action principle
Inspired by the presentation given in Souriau [1964], we start from the following principles
@@
    called presence of the phenomenon in Souriau [1964].
@@
    is stationary for every variation of $G,A^{*}$ and the concomitant phenomena, null on the boundary of an open domain $\mathcal{D}$ of the space-time, $vol$ being the Riemannian volume 4-form.
As the signature of Minkowski metric is (+)(+---)( + - - - ), $\det G<0$, then the component of $vol$ is
@@
Denoting $d$ the exterior derivative of a differential form $\iota_{\bm{V}}$ and the interior product with the vector $\bm{V}$, owing to Cartan’s formula and Stokes’ theorem
@@ 9.2 Laws of conservation
[…]
@@ 9.3 Laws of conservation of the matter
In Souriau [1964, §36], the motion of a continuum is described by a line bundle $\pi_{A}:\mathcal{U}\rightarrow\mathcal{M}$ where $\mathcal{M}$ is a Riemannian manifold of dimension 3 representing the matter and called the matter manifold. Its metric is denoted $\bm{G}_{m}$. Each fiber is the trajectory of a material particle $\bm{a}=\pi_{A}(\bm{X})$. In a local chart of $\mathcal{M}$, it is represented by its Lagrange coordinates $a\in\mathbb{R}^{3}$. As $a$ is an invariant of the motion, the material derivative vanishes:
@@
The vector $\hat{U}$ tangent to the trajectory of a particle $\hat{\bm{a}}$ is such that
@@
where $\Phi$ is a 4-row and $\mu$ is a non null scalar (for $\hat{M}$ being of full rank). If $\mu=0$, A non null solution is $U=0,\hat{U}^{5}\neq 0$, then $\hat{U}$ is not timelike. In the sequel, we exclude this case.
@@
in which $G^{\prime}_{0}$ is Minkowski’s metric (69) and $G_{m}$ is Gram’s matrix of $\bm{G}_{m}$. As in Souriau [1964], we suppose that the behavior of the matter is isotropic
@@
The matrix $H$ represents in a local frame of $\mathcal{M}$ a skew-symmetric tensor
@@
Now we follow the construction proposed in Souriau [1964, (36.7)]. Let $\bm{a}\mapsto r_{0}(\bm{a})$ be a non-vanishing scalar field on $\mathcal{M}$ representing the density of an extensive physical quantity (for instance, the mass or the electric charge), $vol$ be the Riemannian volume 4-form on the space-time $\mathcal{U}$, $vol_{m}$ be the Riemannian volume 3-form on the matter manifold $\mathcal{M}$, and $vol_{r_{0}}$ be the volume 3-form $r_{0}\,vol_{m}$. We define on the space-time the vector field $\bm{J}_{r}$ such that
@@
In Souriau [1965, Chapter 5, §14, Theorem (16)], it is proved that this vector exists and is unique. As $r_{0}\neq 0$, (118) shows that $\bm{J}_{r}$ vanishes only if the rank of $M$ is less than 3. Otherwise, in choosing $Md_{1}X=J_{r}$, we see that
@@
then $\bm{J}_{r}(\bm{X})$ is tangent to the trajectory passing by $\bm{X}$. In Souriau [1964, (36.13)], it is proved that
@@
On this ground, our aim is to construct a Lagrangian $L_{M}$ for the matter that is invariant for every diffeomorphism $\varphi$ of the space-time $\mathcal{U}$
@@
$L_{e}(a,M,A^{*})=\rho_{e0}(a)\,\det\left[\begin{array}[]{c}M\hfill\\ A^{*}\hfill\\ \end{array}\right]=-A^{*}J_{\rho_{e}}=-\rho_{e}\,A^{*}U$
@@ 9.4 Field equations of the gravitation and electromagnetism
According to Souriau [1964, Chapter 5, §14, (35.8) to (35.10)], the Lagrangian of the geometry depends on the connection and its partial derivatives through the curvature tensor
@@
$L_{g}(\Gamma,\bm{R},\bm{R})=-\Lambda+\frac{1}{2}\,G^{ij}R_{ij}$
@@
We recover Einstein’s equations for the gravitation. Einstein’s constant (with $c=1$ and the gravitational constant $G_{N}$ with the index $N$ in Newton’s honour)
@@
Our aim now is to discuss the classical limits of these field equations, starting with (134). First, we consider the first term. In the Galilean approximation (Duval et al. [1985, 1991], Souriau [1997a], de Saxcé and Vallée [2016]), the connection is associated to Galilei group and Christoffel’s symbols $\Gamma^{m}_{ij}$ vanish except777with the convention that the index $t$ corresponds to time and $a,b,c,d$ to the spatial coordinates.
@@
$-\tilde{k}\,\partial_{j}F^{jt}=\kappa\,\rho_{e}U^{i}$
@@
Remark. Arguably, this particularly very small coupling term in (133) can be related to Dirac’s Large Number Hypothesis (LNH) that the ratio of the electrical to the gravitational forces between a proton and an electron of electric charge $e$
@@
where, according to Souriau [1964, Chapter 7, §42, p. 412]
$l_{K}=8\,\pi^{3/2}\,\frac{h\,\sqrt{G_{N}\epsilon_{0}}}{e\,c}=0.238\times 10^{-30}\,\mbox{cm}$
@@ 10 Conclusion
The soon idea of a fifth dimension, if used in a suitable geometric formalism taking into account its very small size, seems now more than ever to offer a promising future for new developments among them we can emphasize:
@@
    For the characterization of charged elementary particles, we largely relied on Souriau [1970, 1997b] without releasing the full potential of the theoretical framework, in particular the construction of the symplectic form and the geometric quantization. These issues have not been considered yet in this work but would deserve to be investigated.
@@
    The next step would be to extend Kaluza-Klein theory to a non-abelian gauge group as in Kerner [1968] with the mathematical tools developed in this paper, in particular the coadjoint orbit method.
@@
    Going back in time, it would worth to revisit works on the early universe cosmology that we think they did not receive the welcome they deserved because based on Kaluza-Klein theory whose weaknesses were known. They could provide a better understanding for Dirac’s large number hypothesis as suggested by Chodos and Detweiler [1980] and to offer a resolution to the horizon problem as claimed by Sahdev [1984].
@@ Acknowledgements
The author would like to thank Leonid Ryvkin for discussions that help to clarify certain points and to Richard Kerner whose advice was very valuable to me. He is also grateful to Claude Vallée for drawing author’s attention on the reference Cartan [1934].

### v2

@@ 8.3 Explicit form of the connection and the equation of motion
Our goal now is to determine explicitly the connection satisfying the hypotheses (H3) to (H5). As we know, the $\hat{\mathbb{G}}_{0}$-structure $\mathcal{F}_{\hat{\bm{G}_{0}}}$ is in general non integrable. Then, instead of the moving frame $(\hat{\bm{e}}^{\prime}_{1},\cdots,\hat{\bm{e}}^{\prime}_{5})$ of (H3), section of $\mathcal{F}_{\hat{\bm{G}_{0}}}$, we would like to work with a coordinate frame $(\bm{\partial}_{1},\cdots,\bm{\partial}_{4})$ associated to a coordinate system $(X^{1},\cdots,X^{4})$ of the space-time $\mathcal{U}$ and the corresponding fibered frame $(\hat{\bm{e}}_{1},\cdots,\hat{\bm{e}}_{5})$ adapted to the section $f$, such as $\hat{\bm{e}}_{i}=(Tf)\bm{\partial}_{i}$ and obtained by the change of basis
@@
As expected, according to the experience, the rest mass $m_{0}$ is an integral of the motion [^5] However, it is possible to extend the formalism to a particle with time-varying mass by introducing a thrust force as in [10]. Besides, (91) is reduced to
@@
represents a 1-covariant and 1-contravariant tensor $\bar{\bm{F}}$ of components $F^{i}_{j}$. The physical interpretation is that the vector $\bm{A}$ of which the adjoint $\bm{A}^{*}$ is represented by the row $A^{*}$ is the electromagnetic 4-potential, $\bm{F}$ (resp. $\bar{\bm{F}}$) is the 2-covariant (resp. 1 contravariant and 1-covariant) electromagnetic field 666with the conventions used in §15 of [48, 50] and
@@ 8.4 Gauge transformation
The coordinate basis $(\bm{\partial}_{i})$ and $\hat{\bm{e}}_{5}$ being given, the adapted basis $(\hat{\bm{e}}_{\alpha})$ of $\mathcal{U}$ to $f$ is uniquely defined by the section $f$. We would like to know to what extend the equation of motion does not depend on the choice of the section $f$. For this aim, we consider another section $f^{\prime\prime}$ and the corresponding adapted basis $(\hat{\bm{e}}^{\prime\prime}_{\alpha})$. As $(T\pi)\hat{\bm{e}}_{i}=(T\pi)\hat{\bm{e}}^{\prime\prime}_{i}=\bm{\partial}_{i}$, the vector $\hat{\bm{e}}^{\prime\prime}_{i}-\hat{\bm{e}}_{i}$ is tangent to the fiber of $\mathcal{U}$ over $\bm{X}$ then
@@ 9 Extended variational relativity
The goal of this section is to extend the general theorems of the relativity based on Einstein-Hilbert functional [21]. We use Palatini formalism, considering the connection as an unknown independent field [38].
@@ 9.1 Stationary action principle
Inspired by the presentation given in [45], we start from the following principles
@@
    called presence of the phenomenon in [45].
@@
    is stationary for every variation of $G,A^{*}$ and the concomitant phenomena, null on the boundary of an open domain $\mathcal{D}$ of the space-time, $vol$ being the Riemannian volume 4-form of component $u$ in the local chart $(X^{i})$
As the signature of Minkowski metric is (+)(+---)( + - - - ), $\det G<0$, then
@@
Denoting $d$ the exterior derivative of a differential form and $\iota_{\bm{V}}$ the interior product with the vector $\bm{V}$, owing to Cartan’s formula and Stokes’ theorem
@@ 9.2 Laws of conservation
The equations $W_{(z)}=0\,$ and $\mbox{div}\,\tilde{\bm{T}}_{(z)}=0\,$ entail the law of conservation
$\mbox{div}\,\bar{\bm{T}}_{(z)}+\tilde{\bm{T}}_{(z)}\cdot\bm{F}=\bm{0}$ (110)
@@ 9.3 Laws of conservation of the matter
In §36 of [45] and §11.2 of [44], the motion of a continuum is described by a line bundle $\pi_{A}:\mathcal{U}\rightarrow\mathcal{M}$ where $\mathcal{M}$ is a Riemannian manifold of dimension 3 representing the matter and called the matter manifold. Its metric is denoted $\bm{G}_{m}$. Each fiber is the trajectory of a material particle $\bm{a}=\pi_{A}(\bm{X})$. In a local chart of $\mathcal{M}$, it is represented by its Lagrange coordinates $a\in\mathbb{R}^{3}$. As $a$ is an invariant of the motion, the material derivative vanishes:
@@
The vector $\hat{\bm{U}}$ tangent to the trajectory of a particle $\hat{\bm{a}}$ is such that
@@
where $\Phi$ is a 4-row and $\mu$ is a non null scalar (for $\hat{M}$ being of full rank). In contrast, if $\mu=0$, a non null solution is $U=0,\hat{U}^{5}\neq 0$, then $\hat{U}$ is not timelike. In the sequel, we exclude this case.
@@
in which $G^{\prime}_{0}$ is Minkowski’s metric (69) and $G_{m}$ is Gram’s matrix of $\bm{G}_{m}$. As in [45], we suppose that the behavior of the matter is isotropic
@@
The matrix $H$ represents in a local frame of $\mathcal{M}$ a self-adjoint tensor
@@
Now we follow the construction proposed in (36.7) of [45]. Let $\bm{a}\mapsto r_{0}(\bm{a})$ be a non-vanishing scalar field on $\mathcal{M}$ representing the density of an extensive physical quantity (for instance, the mass or the electric charge), $vol$ be the Riemannian volume 4-form on the space-time $\mathcal{U}$, $vol_{m}$ be the Riemannian volume 3-form on the matter manifold $\mathcal{M}$, and $vol_{r_{0}}$ be the volume 3-form $r_{0}\,vol_{m}$. We define on the space-time the vector field $\bm{J}_{r}$ such that
@@
In Chapter 5, §14, Theorem (16) of [46], it is proved that this vector exists and is unique. As $r_{0}\neq 0$, (117) shows that $\bm{J}_{r}$ vanishes only if the rank of $M$ is less than 3. Otherwise, in choosing $Md_{1}X=J_{r}$, we see that
@@
then $\bm{J}_{r}(\bm{X})$ is tangent to the trajectory passing by $\bm{X}$. In (36.13) of [45], it is proved that
@@
On this ground, our aim is to construct a Lagrangian $L_{M}$ for the matter that is invariant by every diffeomorphism $\varphi$ of the space-time $\mathcal{U}$
@@
$L_{e}(a,M,G,A^{*})=\rho_{e0}(a)\,\det\left[\begin{array}[]{c}M\hfill\\ A^{*}\hfill\\ \end{array}\right]=-A^{*}J_{\rho_{e}}=-\rho_{e}\,A^{*}U$
@@ 9.4 Field equations of the gravitation and electromagnetism
According to Chapter 5, §14, (35.8) to (35.10) of [45], the Lagrangian of the geometry depends on the connection and its partial derivatives through the curvature tensor
@@
$L_{g}(\Gamma,\bm{R},\bm{G})=-\Lambda+\frac{1}{2}\,G^{ij}R_{ij}$
@@
We recover Einstein’s equations for the gravitation. Einstein’s constant (where $c=1$ and the gravitational constant $G_{N}$ with the index $N$ in Newton’s honour)
@@
Our aim now is to discuss the classical limits of these field equations, starting with (133). First, we consider the former term. In the Galilean approximation ([12, 13, 49, 10]), the connection is associated to Galilei group and Christoffel’s symbols $\Gamma^{m}_{ij}$ vanish except777with the convention that the index $t$ corresponds to time and $a,b,c,d$ to the spatial coordinates.
@@
$-\tilde{k}\,\partial_{j}F^{jt}=\kappa\,\rho_{e}$
@@
Remark. Arguably, this particularly very weak coupling term in (132) can be related to Dirac’s Large Number Hypothesis (LNH) that the ratio of the electrical to the gravitational forces between a proton and an electron of electric charge $e$
@@
where, according to Chapter 7, §42, p. 412 of [45], in SI units
$l_{K}=8\,\pi^{3/2}\,\frac{h\,\sqrt{G_{N}\epsilon_{0}}}{e\,c}=2.38\times 10^{-31}\,\mbox{cm}$
@@ 10 Conclusion
The sound idea of a fifth dimension, if used in a suitable geometric formalism taking into account its overwhelmingly small size, seems now more than ever to offer a promising future for new developments among them we can emphasize:
@@
    For the characterization of charged elementary particles, we largely relied on [48, 50] without releasing the full potential of the theoretical framework, in particular the geometric quantization. These issues have not been considered yet in this work but would deserve to be investigated.
@@
    The next step would be to extend Kaluza-Klein theory to a non-abelian gauge group as in [24] with the mathematical tools developed in this paper, in particular the coadjoint orbit method.
@@
    Going back in time, it would worth to revisit works on the early universe cosmology that we think they did not receive the welcome they deserved because based on Kaluza-Klein theory whose weaknesses were known. They could provide a better understanding for Dirac’s large number hypothesis as suggested in [7] and to offer a resolution to the horizon problem as claimed in [40].
@@ Acknowledgements
The author would like to thank Leonid Ryvkin for discussions that helped to clarify certain points and to Richard Kerner whose advice was very valuable to me. He is also grateful to Claude Vallée for drawing author’s attention on the reference [6].